In 1992, the FAA conducted 86,991 pre-employment drug tests on job applicants who were to be engaged in safety and security-related jobs, and found that 1,143 were positive. (a) Construct a 95 percent confidence interval for the population proportion of positive drug tests
Answer by stanbon(75887)

In 1992, the FAA conducted 86,991 pre-employment drug tests on job applicants who were to be engaged in safety and security-related jobs, and found that 1,143 were positive.
-----------------
p-hat = 1143/86991 = 0.013
---------------------
(a) Construct a 95 percent confidence interval for the population proportion of positive drug tests
---
SE = 1.96*sqrt[0.013*0.987/86991] = 0.000753
===============
95% CI: 0.013-0.000753 < p < 0.013+0.000753
------
Cheers,
Stan H.